OpenStudy (anonymous):

Find the exact value by using a half-angle identity. sin (7pi/8)

OpenStudy (anonymous):

i need answers in radicals:)

OpenStudy (aravindg):

express 7pi/8 as sum or difference of two angles

OpenStudy (anonymous):

can you explain further....?

OpenStudy (e.cociuba):

sin(7pi/8) = +_ ( sqrt( 1- cos(14pu/8)/2) = sqrt ( (1- cos7pi/4)/2 ) = sqrt ( 1-sqrt(3)/2 )/2 ) =sqrt (2 -sqrt(3))/2 It depends on the angel and assumption that you will take positive or negative

OpenStudy (anonymous):

\[ \sin^2(x/2) = \frac{1-\cos(x)}{2} \implies |\sin(x/2)| = \sqrt{\frac{1-\cos(x)}{2}} \]

OpenStudy (anonymous):

\[ x/2 = 7\pi/8 \implies x=7\pi/4 \]

OpenStudy (anonymous):

Since \(0 \lt x/2\lt \pi\) we know that \(\sin\) will be positive.

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

It's expected that you know \(\cos(7\pi/4)\) because it is a simple angle.

OpenStudy (anonymous):

yeaaaa

OpenStudy (anonymous):



OpenStudy (anonymous):

yes...what do i after?

OpenStudy (anonymous):

Okay did you find out \(\cos(7\pi/4)\)?

OpenStudy (anonymous):

yes! 1/sqrt2? @wio

OpenStudy (anonymous):

It's negative.

OpenStudy (anonymous):

oh.:(

OpenStudy (anonymous):

Use this formula: \[ \sin(7\pi/8) = \sqrt{\frac{1-\cos(7\pi/4)}{2}} \]

OpenStudy (anonymous):

@wio okay so would that be my answer when i solve for it??

OpenStudy (anonymous):

@wio i got -1/2sqrt(1-sqrt2) is this right?

OpenStudy (anonymous):

Umm, not sure how you got that.

OpenStudy (anonymous):

@wio hmmmm -1/2sqrt(2-sqrt2) maybe that's it?

OpenStudy (anonymous):

like this??